OpenStudy (anonymous):

Convert r = 4sin theta sec ^{2} theta to a rectangular equation

OpenStudy (freckles):

I see no equation...

OpenStudy (freckles):

Is there suppose to be one?

OpenStudy (freckles):

\[r \sin(\theta) \sec^2(\theta) \text{ is a polar expression }\]

OpenStudy (freckles):

\[\text{ use the following } \\ r^2=x^2+y^2 \\ r \sin(\theta)=y \\ r \cos(\theta)=x \\ \tan(\theta)=\frac{y}{x}\]

OpenStudy (anonymous):

\[r = 4\sin \theta \sec ^{2} \theta \] is the correct equation... i am so sorry

OpenStudy (freckles):

hmm... maybe we can try to multiply r on both sides as a first step

OpenStudy (freckles):

\[r^2=4 \cdot r \sin(\theta) \cdot \sec^2(\theta)\] so we know we can replace r^2 with x^2+y^2 and rsin(theta) with y

OpenStudy (freckles):

then try to divide sec^2(theta) on both sides

OpenStudy (freckles):

and see if you can figure it out from there i think it shouldn't look to bad :)

OpenStudy (anonymous):

alright, give me one second to write it out!

OpenStudy (anonymous):

okay, but what do i do with that sec^2 theta on the bottom?

OpenStudy (freckles):

1/sec^2(theta)=cos^2(theta)

OpenStudy (freckles):

\[r^2 \cos^2(\theta)=4 \cdot r \sin(\theta)\]

OpenStudy (freckles):

if x=rcos(theta) then x^2=?

OpenStudy (anonymous):

x^2 = rcos(theta)^2 ?

OpenStudy (freckles):

\[x^2=(r \cos(\theta)))^2\] \[x^2=r^2 \cos^2(\theta)\]

OpenStudy (freckles):

and that is exactly what you have on the left of your equation

OpenStudy (anonymous):

okay so dont change the r^2 to x^2 + y^2 then?

OpenStudy (freckles):

nope we are going to use that r^2 with the cos^2(theta) to right the left hand side has x^2

OpenStudy (freckles):

to write*

OpenStudy (anonymous):

alrighty

OpenStudy (freckles):

and you know to replace the rsin(theta) with?

OpenStudy (anonymous):

y ?

OpenStudy (freckles):

yep yep

OpenStudy (anonymous):

okay so the ending converted equation is something like \[x ^{2} - 4y\]

OpenStudy (freckles):

x^2=4y

OpenStudy (freckles):

and i think that is what you meant and yes

OpenStudy (anonymous):

oh okay, thank you so much

OpenStudy (freckles):

I hope everything made sense :)

OpenStudy (anonymous):

yes! your help is very much appreciated

OpenStudy (freckles):

thanks